Mariam just inherited an amount of money. While she plans to squander some of it away, how much should she deposit in an account earning 5% interest per year if she would like to have OMR 10,000 in the account in 10 years?
1
Expert's answer
2020-11-16T06:51:35-0500

In this case we use the following formula

where P.V is present value

F.V is future value

10000 = "P.V\times [(1+0.05)^{10} ]"


"F.V = P.V \times [(1+i)]^n"

"10000 = P.V\times[(1+0.05)]^{10}"

10000=1.62889PV

"P.V= \frac{1000}{1.05^{10}}"


P.V =6139.13